# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v2

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast. To this end, several learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach for guided reconstruction addressing this issue, which additionally provides an explanatory framework for the multi-contrast problem in terms of the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased content of the image estimate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-MUNIT. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house multi-coil raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
I Introduction
@@
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4, 5] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [6].
@@
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [7, 8, 9] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [10, 11, 12]. In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data. The two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. This illustration only provides an intuition about the information encoded in the content and style domains, and does not represent the actual training process, which is given in detail in Section III-B. (b) The reconstruction stage applies this model as a content consistency operator (bottom) within an ISTA-based iterative algorithm. Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference. CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator. DC denotes data consistency. Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS-WT and PnP-CNN reconstruction, respectively, are shown for comparison.
@@
The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS [7, 8] to end-to-end learning with unrolled networks [9, 13, 14, 15] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [16]. Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data. We address this issue by proposing a plug-and-play reconstruction method that splits the problem into a purely image-domain learning sub-problem and an iterative reconstruction sub-problem. The learning problem leverages ideas from content/style decomposition, thereby offering a degree of interpretability to our approach. To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
@@
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [17, 18, 19, 20]. While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast. Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy. That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [21, 22], which can be adopted to complement image reconstruction. We observe that methods such as MUNIT [22] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
@@
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line [23, 24] has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
@@
We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This training process is independent of the reconstruction problem and can be performed using unpaired images. We then make the interesting observation that in multi-contrast MR images, the style information tends to localize in the center of the k-space, allowing for an accurate style estimation from an undersampled image. Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step. We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-MUNIT [25]. An overview of our approach is shown in Fig. 1. While the PnP-based decoupling of image modeling and image reconstruction simplifies the training process and allows the two stages to be analyzed separately, a further level of decoupling offered by content/style disentanglement offers additional modularity and an intuitive guidance mechanism for the reconstruction.
Specifically, our contributions are four-fold:
@@
    Developing this idea further and incorporating a corrective process for the content, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
@@
    Through comprehensive experiments, we shed light on several properties of PnP-MUNIT such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
II Related Work
II-A Reconstruction Methods for Accelerated MRI
@@
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) [3], which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA [26] or ADMM [27] family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods [23, 24] replace the proximal operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks [28, 4, 5] extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability [23].
@@
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [7] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. [8] proposed adaptive weighting-based guided CS and Song et al. [29] used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods [9, 13, 30, 31, 32, 15], on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet [15] is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
@@
II-B Unpaired Image-to-Image Modeling
@@
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [33] and CycleGAN [34] in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT [21], assumes a shared latent space underlying the two domains to explicitly represent shared information. However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [17, 35, 18]. However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. [19] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
@@
MUNIT [22] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as [36, 37, 38], MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to [19], we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model these using MUNIT, referring to the shared and non-shared components as content and style, respectively.
II-C Combining Image Translation with Reconstruction

### v3

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast. To this end, several end-to-end learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach addressing this issue, additionally providing an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-CoSMo. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a small radiological task, PnP-CoSMo allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
keywords:
MRI reconstruction , multi-modal , content/style decomposition , plug-and-play
journal: Medical Image Analysis
[…]
1 Introduction
@@
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging (Pruessmann et al., 2001; Griswold et al., 2002), compressed sensing (CS) (Lustig et al., 2007), and deep learning reconstruction (Aggarwal et al., 2018; Pezzotti et al., 2020) have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow (Seow et al., 2024).
@@
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan (Ehrhardt and Betcke, 2016; Weizman et al., 2016; Zhou and Zhou, 2020) and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously (Bilgic et al., 2011; Huang et al., 2014; Kopanoglu et al., 2020). In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data. We assume that the two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ can be decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. This model is learned in two stages – an unpaired pre-training stage and a paired fine-tuning stage, both requiring only image data. (b) The reconstruction stage applies the content/style model as a content consistency operator (bottom) within an ISTA-based iterative scheme. Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference. The ”refine” block denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator. DC denotes data consistency. For comparison, two other reconstruction priors are shown, namely wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS and PnP-CNN algorithms, respectively.
@@
The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS (Ehrhardt and Betcke, 2016; Weizman et al., 2016) to end-to-end learning with unrolled networks (Zhou and Zhou, 2020; Yang et al., 2020; Pooja et al., 2022; Lei et al., 2023) and, more recently, diffusion model-based Bayesian maximum a posteriori estimation (Levac et al., 2023). Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data. We address this issue by proposing a plug-and-play reconstruction method that splits the problem into a purely image-domain learning sub-problem and an iterative reconstruction sub-problem. The learning problem leverages ideas from content/style decomposition, thereby offering a degree of interpretability to our approach. To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
@@
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another (Dar et al., 2019; Yurt et al., 2021; Denck et al., 2021; Atli et al., 2024). While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast. Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy. That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data (Liu et al., 2017; Huang et al., 2018), which can be adopted to complement image reconstruction. We observe that methods such as MUNIT (Huang et al., 2018) can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
@@
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line (Ahmad et al., 2020; Kamilov et al., 2023) has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
@@
We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This training process is independent of the reconstruction problem and can be performed using unpaired images. We then make the interesting observation that in multi-contrast MR images, the style information tends to localize in the center of the k-space, allowing for an accurate style estimation from an undersampled image. Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step. We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-CoSMo (Plug-and-Play method based on Content/Style Modeling) (Rao et al., 2024). An overview of our approach is shown in Fig. 1. While the PnP-based decoupling of image modeling and image reconstruction simplifies the training process and allows the two stages to be analyzed separately, a further level of decoupling offered by content/style disentanglement offers additional modularity and an intuitive guidance mechanism for the reconstruction. Specifically, our contributions are four-fold:
@@
    Developing this idea further and incorporating a corrective process for the content, we propose PnP-CoSMo, a modular algorithm for guided reconstruction combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
@@
    Through comprehensive experiments, we shed light on several properties of PnP-CoSMo such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility with a specific radiological task.
2 Related work
2.1 Reconstruction methods for accelerated MRI
@@
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) (Lustig et al., 2007), which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA (Chambolle et al., 1998) or ADMM ((Boyd et al., 2011)) family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods (Ahmad et al., 2020; Kamilov et al., 2023) replace the citep operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks (Schlemper et al., 2017; Aggarwal et al., 2018; Pezzotti et al., 2020) extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability (Ahmad et al., 2020).
@@
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke (2016) introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. (2016) proposed adaptive weighting-based guided CS and Pingfan et al. (2018) used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods (Zhou and Zhou, 2020; Yang et al., 2020; Liu et al., 2021a, b; Feng et al., 2022; Lei et al., 2023), on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet (Lei et al., 2023) is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
@@
2.2 Unpaired image-to-image modeling
@@
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix (Isola et al., 2017) and CycleGAN (Zhu et al., 2017) in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT (Liu et al., 2017), assumes a shared latent space underlying the two domains to explicitly represent shared information. However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis (Dar et al., 2019; Oh et al., 2020; Yurt et al., 2021). However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. (2021) partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
@@
MUNIT (Huang et al., 2018) extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same ”content” but differing in ”style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as (Kotovenko et al., 2019; Kwon and Ye, 2021; Wu et al., 2023), MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to Denck et al. (2021), we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model using an approach based on the MUNIT framework, referring to the shared and non-shared components as content and style, respectively.
2.3 Combining image translation with reconstruction